Summary: Heavy cloud cover at the beginning, but cleared up enough to start observing at 3:40 (but had some clouds most of the time). Dome again closed for 2.5 hrs due to heavy clouds & high humidity. Observing was supported by Dave Thompson by phone because Olga had a family emergency. Ian McGreer later joined by videocon. Got confused by the fact that Lucifer slitmasks do not have alignment holes physically cut out. This was actually fine because Lucifer's mask alignment procedure doesn't require imaging alignment stars through a slitmask (but is it really a good idea not to have alignment holes on slitmasks?). Obtained only four 10min spectra for Egami's program (partially through clouds). Clearly saw signal from a few bright sources, confirming that the MOS mode is working fine. Without a support person on-site, the whole operation was not as straightforward as it could have been (especially when we encountered various problems), but thanks to Dave and Ian's help, we did manage to execute some MOS observations.


  • Taking flats and arcs while waiting for clouds to clear up.
3:40
  • Dome opened.
    Had a problem starting up LUCIFER in LBCTools. Dave had to modify the program.
4:00
  • Pointing check.
    High-elevation star - ptcorr; Low-elevation star - manual IA offset; High-elevation star - ptcorr
4:30
  • Slewing to a Persson standard close to the first target (A1703).
    Collimation.
4:50
  • Slewing to the first target (A1703).
    Taking mask image.
4:50-5:30
  • Was going to perform mask alignment by ourselves while waiting for Dave Thompson to drive to Steward from home.
    However, got confused by the fact that there were no alignment holes on the mask.
    Checking the submitted mask design files, etc.
    Eventually learned from Dave after his arrival at Steward that this is fine. Lucifer mask alignment procedure doesn't require imaging alignment stars through alignment holes.
5:30
  • Mask alignment procedure started.
    Taking mask image.
    Taking sky images.
    Aligning the mask and adjusting the telescope pointing.
    Taking mask image and making sure that bright sources are in the slitlets.
6:10
  • Starting MOS integration for Egami's program (150-Ks).
    Lost the guide star after two 10min integrations due to clouds.
    Humidity shooting up. Heavy cloud cover.
6:30
  • Dome closed....
9:00
  • Dome reopened.
9:20
  • Couldn't move masks; initializing the Instrument Control GUI. %% Restarted the GUI after calling up Dave at home.
9:40
  • Slewed to a Persson star close to A. Zabludoff's target.
    Pointing check; Collimation. Having trouble collimating.
    Running out of time on this target.
10:05
  • Moving to G. Rieke's 1st target.
    Guide star not found for the 1st target.
10:20
  • Moving to G. Rieke's 2nd target.
    Seeing ~1.2-1.4" Target too faint; couldn't acquire on the slit.
10:35
  • Switching back to Egami's target.
    Pointing check; Collimation.
11:05
  • Starting MOS integration.
    Taking 2 10min integrations.
11:30
  • Slewing to a Telluric star.
    Guide star not found. Trying to acquire the star in the tracking mode, but had trouble making precise offsets.
12:15
  • Dome closed.

-- EiichiEgami - 17 Apr 2010
Topic revision: r4 - 18 Apr 2010, XiaohuiFan
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback